OpenStudy (anonymous):

give the vaules of all the vertical asymptotes,horizonal asympyotes,slant asymptotes, and the holes to the equation (2x-6) / (x^2 - 3x)

OpenStudy (amistre64):

factor the top and bottom so that you can cancel out like parts

OpenStudy (anonymous):

how do u do that i do not understand it period

OpenStudy (amistre64):

if you cannot factor a polynomial yet, then you are not ready to do this kind of problem

OpenStudy (anonymous):

i have to its my class work in percal

OpenStudy (amistre64):

then you have to learn how to factor polynomials first, or you will not be able to work out the maths the build on it.

OpenStudy (amistre64):

2x-6 do 2x and 6 have any common factors?

OpenStudy (anonymous):

yes 12

OpenStudy (amistre64):

12 would be a common multiple; not a common factor the factors of: 2x: 2,x 6: 2,3 notice that the common factor that they share is "2"

OpenStudy (anonymous):

wats the difference between a common factor and a common multiple and yes i see that

OpenStudy (amistre64):

factors are the smaller parts that make up a number; a multiple is what you get when you multiply the numbers together. they are like 2 sides of the same coin. different, but related.

OpenStudy (anonymous):

oh okay

OpenStudy (amistre64):

\[2x-6 = 2(x) - 2(3)\] \[2(x) - 2(3)=2(x-3)\]this is the "factored" form for the top

OpenStudy (amistre64):

lets factor the bottom x^2 - 3x the factors of: x^2: x, x 3x: 3, x notice that the factor that they have in common is "x"

OpenStudy (anonymous):

I recommend a quick refresher when you get a chance, @lissia_liverman http://themathpage.com/aPreCalc/algebraPre.htm

OpenStudy (anonymous):

thanks CliffSedge

OpenStudy (amistre64):

\[x^2 - 3x=x(x)-x(3)\] \[x(x)-x(3)=x(x-3)\]is the factored form of the bottom, giving us \[\frac{2x-3}{x^2-3x}~=~\frac{2(x-3)}{x(x-3)}\]

OpenStudy (amistre64):

with this in the factored form, we can start to recognize the holes and vertical asympotes

OpenStudy (anonymous):

u no ur good at explaining things amistre64

OpenStudy (amistre64):

lol, im just too old to know better :)

OpenStudy (anonymous):

lol nice i like that :)

OpenStudy (amistre64):

"holes" are what happens when we can remove things. notice that the like factors top to bottom can be "removed"\[\frac{2\cancel{(x-3)}}{x\cancel{(x-3)}}\] therefore when x-3 = 0 we have a "hole" the vertical asymptotes are what we are left with that zero out the bottom \[\frac{2}{x} \]in this case, when x=0 we get a zero in the bottom and a vertical asymptote

OpenStudy (amistre64):

the horizontal and slant asymptotes are what happens when the values of x get very big or very small.

OpenStudy (amistre64):

\[\frac{2x-6}{x^2 - 3x}\]when x get very big (or very small) the leading terms take control of the setup. for example:\[\frac{2(1,000,000)-6}{(1,000,000)^2-3(1,000,000)}\]is "close to" \[\frac{2(1,000,000)}{(1,000,000)^2}\] the other parts dont contribute enough to the setup to matter at these large values

OpenStudy (anonymous):

oh okay how would you graph this original problem

OpenStudy (amistre64):

so, to find the horizontal (or slant) we focus on simplifying the leading terms only and see how they act when x grows into infinity \[\frac{2x...}{x^2...}\to\ \frac{2}{x}\] as x gets larger and larger, this setup gets closer and closer to zero. Which will be the horizontal asymptote then graphing it? that might be a bit complicated by hand.

OpenStudy (amistre64):

i can determine a sketch of it using the information from the stuff we have already found to make a close approximation to how it should look

OpenStudy (anonymous):

thank u

OpenStudy (amistre64):

 and then to fill in the way it cozies up to the VA and the HA

OpenStudy (amistre64):

\[\frac{2(x-3)}{x(x-3)}\] lets test a value of x that is clearly between 0 and 3, say x=1. And we are interested in the sign of the answer rather than the value itself, the sign tells us what side of the x axis we want to be on \[\frac{2(1-3)}{1(1-3)}\] \[\frac{+(-)}{+(-)}=+\] everything between 0 and 3 will be positive; lets use x=4 since it is on the other side of the 3 \[\frac{2(4-3)}{4(4-3)}\] \[\frac{+(+)}{+(+)}=+\] And x=-4 is good enough for the other side of 0 \[\frac{2(-4-3)}{-4(-4-3)}\] \[\frac{+(-)}{-(-)}=-\]  so something similar to this